OpenStudy (anonymous):

Triangle ABC is similar to Triangle XYZ. What is the length of segment XZ if BC=9m,AC=6m and YZ=3m?

hero (hero):

Hints: 1. In general, when two triangles are similar, the ratio of two sides of one triangle, equals the ratio of the corresponding sides of the other triangle 2. In this case, \(\frac{AC}{BC} = \frac{XZ}{YZ}\), so in other words, \(\frac{6}{9} = \frac{x}{3}\) 3. Solve for x to get the length of XZ

OpenStudy (anonymous):

Huh ? sorry i dont really get it x)

hero (hero):

All you have to do is solve the proportion for x. You do know how to solve proportions, don't you?